# Unified Framework of Forced Magnetic Reconnection and Alfvén Resonance

### v1

@@ VI.2 Transition to Alfvén Resonance
In the previous subsection, we analyzed the constant-psi regime of Eq. (28) for small frequencies ($\hat{\omega}_{0}\ll S^{-1/3}$). As discussed, when $\hat{\omega}_{0}\ll S^{-1/3}$ the solution corresponding to the constant-psi regime matches very well with the simulation results. However, as the frequency is increased beyond $\hat{\omega}_{0}\sim S^{-1/3}$, the agreement diverges and the constant-psi solution breaks down. This is shown for example by the dotted blue line in Figure 3, that represents the constant-psi solution but for a high frequency case, $\hat{\omega}_{0}=0.1$; as can be seen, the dotted line does not reproduce simulation results, which are represented by the solid blue line. Therefore, for large frequencies we must consider a different limit to describe Alfvén resonance.
For large frequencies, $\hat{\omega}_{0}\gg S^{-1/3}$, the residues at the resonant poles of the boundary driver, represented in green in Figure 2, contribute to the time-asymptotic response of the system. The contributions from both the jump at the branch cut and the residues at the poles located in the region $\Re({\hat{g}})<0$ have vanishing contributions for $\hat{t}\rightarrow\infty$. Additionally, if the frequency is large enough such that $\hat{\omega}_{0}\gg S^{-1/3}\hat{k}^{2/3}$, one cannot use the asymptotic expansion of Eq. (28) and (30) for $Q\ll 1$ to evaluate the residues at $\hat{g}=\pm i\hat{\omega}_{0}$, as it was done for small frequency. Instead, one should use the asymptotic expansion for $Q\gg 1$ to calculate the residues at those poles, except this time by retaining resistive corrections. As gradients become strong around the resonant point, the effects of resistivity will be to resolve and smooth out the logarithmic singularity of the ideal solution. Performing the same expansion for $Q\gg 1$ as with the ideal case but by retaining higher-order terms in $z/Q^{3/2}\ll 1$, we find that the stream function can be approximated by
$\tilde{\phi}=\frac{\hat{\psi}_{0}}{2}\frac{1}{\hat{\theta}}e^{-\frac{\hat{g}}{|\hat{k}|}\hat{\theta}\left(1+\frac{1}{6}\frac{\hat{\theta}^{2}}{\hat{g}S}\right)+\mathcal{O}\left(\frac{\hat{\theta}^{5}}{\hat{g}S^{2}}\right)}\left[1+\mathcal{O}\left(\frac{\hat{\theta}^{2}}{\hat{g}S}\right)\right],$ (51)
@@
$\tilde{\phi}=i\frac{\Xi_{0}}{\pi}\frac{\hat{\omega}_{0}\sin(\hat{k}_{0}\hat{y})}{\sinh(\hat{k}_{0})}\sin\left(\hat{\omega}_{0}(\hat{t}-\frac{\hat{\theta}}{\hat{k}_{0}})\right)\frac{e^{-\frac{1}{6}\frac{\hat{\theta}^{3}}{S\hat{k}_{0}}}}{\hat{\theta}}.$ (52)
To find the spatio-temporal stream function we finally used Mathematica to evaluate the inverse Fourier transform from $\hat{\theta}$ to $\hat{x}$ of Eq. (52). The resulting solution for $\hat{\omega}_{0}=0.1$ and $\hat{\omega}_{0}=0.05$ is reported in Figure 3, dashed blue and red lines, and compared with simulation results represented by the solid blue and red lines. It is apparent that Eq. (51), when Fourier-inverted, captures the long-time behavior of Alfvén resonances. We see that the contribution of the resistivity is both to remove the logarithmic singularity and to provide a finite width to the resonant boundary layer. The boundary layer, as can be seen by inspection of Eq. (52), scales as $\delta\sim S^{-1/3}$.
@@
$\tilde{J}_{z}=\frac{\hat{\psi}_{0}}{2}\frac{\hat{k}}{\hat{g}}\left(\frac{\hat{g}}{|\hat{k}|}\hat{\theta}+1\right)e^{-\frac{\hat{g}}{|\hat{k}|}\hat{\theta}\left(1+\frac{1}{6}\frac{\hat{\theta}^{2}}{\hat{g}S}\right)}.$ (53)
Again we evaluate the residues at the resonant poles of the boundary condition to find the long time behavior, yielding the following solution to be numerically inverted from $\hat{\theta}$ to configuration space,
$\tilde{J}_{z}=-\frac{\Xi_{0}\hat{k}_{0}}{\pi}\frac{e^{-\frac{1}{6}\frac{\hat{\theta}^{3}}{S\hat{k}_{0}}}}{\sinh(\hat{k}_{0})}\left(\cos(\zeta)-\sin(\zeta)\frac{\hat{\omega}_{0}}{\hat{k}_{0}}\hat{\theta}\right),$ (54)
where $\zeta=\hat{\omega}_{0}(\hat{t}-\frac{\hat{\theta}}{\hat{k}_{0}})$. The current density at the resonant site, found through numerical inversion of (54), has been plotted in the right frame of Figure 4. Just like with the reconnection rate for small frequencies, we see very good agreement between the ideal solution and the numerical simulation up to nearly $\hat{t}\sim S^{1/3}$. Over longer time scales, the system reaches a steady-state due to resistive dissipation, and the solution in Eq. (54) provides the correct time-asymptotic behavior where an oscillating steady-state is reached.
Figures 3 and 4 together demonstrate the transition from Alfvén resonance to forced magnetic reconnection as $\hat{\omega}_{0}$ varies from $\hat{\omega}_{0}\gg S^{1/3}$ to $\hat{\omega}_{0}\ll S^{1/3}$. As the frequency approaches zero, the resonant sites shifts towards the neutral line, since Alfvén resonance necessarily occurs where the phase velocity of the driver is equal to the local Alfvén speed, $\hat{\omega}_{0}/\hat{k}=v_{a}(x)$. When the resonant layers, which are located on either side of the neutral line, overlap across the neutral line as a consequence of their finite widths, the resonance couples with the reconnection mode. The subsequent section discusses the effect of finite frequency on such a coupling.
@@ VI.3 Discussion: scalings laws and effect of finite frequency on reconnection
Figure 5: Plots of the reconnection rate and reconnected flux as functions of $\hat{\omega}_{0}$ evaluated at the time of maximum reconnection rate in the constant-psi regime, for $S=10^{8}$. They are calculated from the inverse Laplace transforms of Eq. (49), divided by $S$, and Eq. (44), which are applicable for $\hat{\omega}_{0}\ll S^{-1/3}$. The red (left most) line in each plot indicates $\hat{\omega}_{0}=S^{-3/5}$. The orange (middle) line indicates the location of half maximum half width of $\psi_{0}$. The green (right most) line indicates $\hat{\omega}_{0}=S^{-1/3}.$ In the frequency range beyond the half-width of $\psi_{0}$ but below $\hat{\omega}_{0}=S^{-1/3}$, we observe a power law scaling $R\sim\hat{\omega}_{0}^{-0.25}$ and $\psi_{0}\sim\hat{\omega}_{0}^{-1.23}$.
@@
To further understand the effect of finite frequency on forced reconnection, in this section we analyze two sets of scalings predicted by the theory. The first set is the scaling of the width $\delta$ of the boundary layer for forced reconnection and Alfvén resonance as functions of the Lundquist number $S$. The second set is the scaling of the constant-psi reconnection rate and reconnected flux at the neutral line as functions of the driving frequency $\hat{\omega}_{0}$.
@@
Through inspection of the current density equation in the constant-psi regime which describes forced reconnection, Eq. (48), we find the theoretical scaling law for $\delta$. For this regime we assume the scaling $\hat{g}\sim S^{-3/5}$, found from the poles in the denominator. Upon substitution of $\hat{g}$ for $S$ in Eq. (48) we find the argument of the Modified Bessel and Struve functions to be of order unity when $\hat{x}\sim S^{-2/5}$. This indicates a layer width scaling $\delta\sim S^{-2/5}$, which is the known scaling for this process [2].
@@
For Alfvén resonance we perform the same analysis of its equivalent current density equation, applicable for large driving frequencies $\hat{\omega}_{0}\gg S^{-1/3}$, Eq. (54). For this scaling we need only the argument of the exponential to determine the inner layer width dependence on $S$. We see that the argument is of order unity when $\hat{\theta}\sim S^{1/3}$ which is indicative of a layer width scaling $\delta\sim S^{-1/3}$, as expected [9, 8, 10].
@@
Although it is not possible to provide a theoretical solution for the intermediate regime when $Q\sim 1$ (which characterizes the region at which our approximations mismatch with simulation in Figure 4), we can nevertheless use the general solution given in Eq. (28) to infer the expected $\delta$. From the definition of the parameter $Q$, Eq. (23), the scaling $\hat{g}\sim S^{-1/3}$ naturally emerges. Replacing $\hat{g}\sim S^{-1/3}$ in Eq. (28) we find the argument of the Tricomi confluent hypergeometric functions of Eq. (28), $z=\sqrt{Q}\hat{\theta}/(S\hat{k})^{2/3}$, is of order 1 for $\hat{\theta}\sim S^{1/3}$ or $\delta\sim S^{-1/3}$. The regime $Q\sim 1$, which corresponds to $\hat{t}\sim S^{1/3}$, applies to both Alfvén resonance and forced reconnection. In particular, it shows that forced reconnection starts off in the non-constant-psi regime, as first discussed in [2], before continuing on through the constant-psi. Thus, we see that when resistivity becomes important at $\hat{t}\sim S^{1/3}$ both forced reconnection and Alfvén resonance exhibit the same layer width, allowing for their coupling, when the forcing frequency is sufficiently small.
@@
We make now analytical predictions of how the reconnection rate and the reconnected flux scale with the forcing frequency $\hat{\omega}_{0}$. We conduct this analysis under the constant-psi regime, i.e., $R$ is calculated by using Eq. (49) divided by $S$, and (44) is used for $\psi_{0}$, that in this approximation represents the reconnected flux at $\hat{x}=0$.
@@
For frequencies $\hat{\omega}_{0}\ll S^{-3/5}$, the system has time to evolve from the ideal to the constant-psi regime almost unaffected by the oscillating driver, and we expect that for such low frequencies both $R$ and $\hat{\psi}_{0}$ are unaffected by $\hat{\omega}_{0}$. When the frequency approaches the inverse of the constant-psi time scale, which is around the time when the peak in reconnection occurs (see Fig. 4), we expect the reconnection process to be affected by the finite frequency driver. Now, in the frequency range $S^{-3/5}\ll\hat{\omega}_{0}\ll S^{-1/3}$ one can neglect the term $\Delta_{0}=2\hat{k}_{0}/\tanh(\hat{k}_{0})$ in the denominator of Eqs. (44) and (49). Additionally, as demonstrated previously, as $\hat{\omega}_{0}$ increases the contribution to the long time solution is provided by residues at $\hat{g}=\pm i\hat{\omega}_{0}$. By analyzing under this light the residues of the inverse Laplace transforms of Eqs. (49) and (44) at $\hat{g}=\pm i\hat{\omega}_{0}$, it can be shown that the expected scalings are $R\sim\hat{\omega}_{0}^{-1/4}$ and $\psi_{0}\sim\hat{\omega}_{0}^{-5/4}$. This would suggest that although $\hat{\omega}_{0}\sim S^{-1/3}$ determines the transition from forced reconnection to Alfvén resonance, the reconnection rate is reduced in the range of frequencies $S^{-3/5}\ll\hat{\omega}_{0}\ll S^{-1/3}$. For frequencies larger than $\hat{\omega}_{0}\gg S^{-1/3}$ reconnection is completely suppressed by the Alfvén resonance.
@@
To measure the effect of finite frequency on reconnection, in Figure 5 we show the plot of the normalized, maximum reconnection rate $R$ and corresponding reconnected flux function $\psi_{0}$ as functions of $\hat{\omega}_{0}$ in the left and right panel, respectively, for $S=10^{8}$. As explained above, the constant-psi regime was assumed in making these plots. Therefore, we used Eqs. (49), divided by $S$, for $R$ and (44) for $\psi_{0}$, and calculated their inverse Laplace transforms at the time for which the reconnection rate is at its maximum. With the vertical red, orange and green (left, central, and right) lines, we mark $\hat{\omega}_{0}=S^{-3/5}$, the half-width of $\psi_{0}$, and $\hat{\omega}_{0}=S^{-1/3}$. As can be seen, the predicted scalings of $R\sim\hat{\omega}_{0}^{-1/4}$ and $\psi_{0}\sim\hat{\omega}_{0}^{-5/4}$ are very closely matched in Figure 5, in which we found power law scalings of $R\sim\hat{\omega}_{0}^{-0.25}$ and $\psi_{0}\sim\hat{\omega}_{0}^{-1.23}$ for the reconnection rate and reconnected flux, respectively, shown by the fits with the dashed line. That is we find, as expected, a strong suppressing effect on reconnection for frequencies less than the transition frequency $\hat{\omega}_{0}\sim S^{-1/3}$ and higher than $\hat{\omega}_{0}\sim S^{-3/5}$.

### v2

@@ VI.2 Transition to Alfvén Resonance
In the previous subsection, we analyzed the constant-psi regime of Eq. (28) for small frequencies ($\omega_{0}\ll S^{-1/3}$). As discussed, when $\omega_{0}\ll S^{-1/3}$ the solution corresponding to the constant-psi regime matches very well with the simulation results. However, as the frequency is increased beyond $\omega_{0}\sim S^{-1/3}$, the agreement diverges and the constant-psi solution breaks down. This is shown for example by the dotted blue line in Figure 3, that represents the constant-psi solution but for a high frequency case, $\omega_{0}=0.1$; as can be seen, the dotted line does not reproduce simulation results, which are represented by the solid blue line. Therefore, for large frequencies we must consider a different limit to describe Alfvén resonance.
For large frequencies, $\omega_{0}\gg S^{-1/3}$, the residues at the resonant poles of the boundary driver, represented in green in Figure 2, contribute to the time-asymptotic response of the system. The contributions from both the jump at the branch cut and the residues at the poles located in the region $\Re({g})<0$ have vanishing contributions for $t\rightarrow\infty$. Additionally, if the frequency is large enough such that $\omega_{0}\gg S^{-1/3}k^{2/3}$, one cannot use the asymptotic expansion of Eq. (28) and (30) for $Q\ll 1$ to evaluate the residues at $g=\pm i\omega_{0}$, as it was done for small frequency. Instead, one should use the asymptotic expansion for $Q\gg 1$ to calculate the residues at those poles, except this time by retaining resistive corrections. As gradients become strong around the resonant point, the effects of resistivity will be to resolve and smooth out the logarithmic singularity of the ideal solution. Performing the same expansion for $Q\gg 1$ as with the ideal case but by retaining higher-order terms in $z/Q^{3/2}\ll 1$, we find that the stream function can be approximated by
$\tilde{\phi}(\theta,y,g)=\frac{\hat{\psi}_{0}}{2}\frac{1}{\theta}e^{-\frac{g}{|k|}\theta\left(1+\frac{1}{6}\frac{\theta^{2}}{gS}\right)+\mathcal{O}\left(\frac{\theta^{5}}{gS^{2}}\right)}\left[1+\mathcal{O}\left(\frac{\theta^{2}}{gS}\right)\right],$ (51)
@@
$\tilde{\phi}(\theta,y,t)=i\frac{\Xi_{0}}{\pi}\frac{\omega_{0}\sin(k_{0}y)}{\sinh(k_{0})}\sin\left(\omega_{0}(t-\frac{\theta}{k_{0}})\right)\frac{e^{-\frac{1}{6}\frac{\theta^{3}}{Sk_{0}}}}{\theta}.$ (52)
To find the spatio-temporal stream function we finally used Mathematica to evaluate the inverse Fourier transform from $\theta$ to $x$ of Eq. (52). The resulting solution for $\omega_{0}=0.1$ and $\omega_{0}=0.05$ is reported in Figure 3, dashed blue and red lines, and compared with simulation results represented by the solid blue and red lines. It is apparent that Eq. (51), when Fourier-inverted, captures the long-time behavior of Alfvén resonances. We see that the contribution of the resistivity is both to remove the logarithmic singularity and to provide a finite width to the resonant boundary layer. The boundary layer, as can be seen by inspection of Eq. (52), scales as $\delta\sim S^{-1/3}$.
@@
$\tilde{J}_{z}(\theta,k,g)=\frac{\psi_{0}}{2}\frac{k}{g}\left(\frac{g}{|k|}\theta+1\right)e^{-\frac{g}{|k|}\theta\left(1+\frac{1}{6}\frac{\theta^{2}}{gS}\right)}.$ (53)
Again we evaluate the residues at the resonant poles of the boundary condition to find the long time behavior, yielding the following solution to be numerically inverted from $\theta$ to configuration space,
$\tilde{J}_{z}(\theta,y,t)=-\frac{\Xi_{0}}{\pi}\frac{k_{0}\cos(k_{0}y)}{\sinh(k_{0})}e^{-\frac{1}{6}\frac{\theta^{3}}{Sk_{0}}}\left(\cos(\zeta)-\sin(\zeta)\frac{\omega_{0}}{k_{0}}\theta\right),$ (54)
where $\zeta=\omega_{0}(t-\frac{\theta}{k_{0}})$. The current density at the resonant site, found through numerical inversion of (54), has been plotted in the right frame of Figure 4. Just like with the reconnection rate for small frequencies, we see very good agreement between the ideal solution and the numerical simulation up to nearly $t\sim S^{1/3}$. Over longer time scales, the system reaches a steady-state due to resistive dissipation, and the solution in Eq. (54) provides the correct time-asymptotic behavior where an oscillating steady-state is reached.
Figures 3 and 4 together demonstrate the transition from Alfvén resonance to forced magnetic reconnection as $\omega_{0}$ varies from $\omega_{0}\gg S^{1/3}$ to $\omega_{0}\ll S^{1/3}$. As the frequency approaches zero, the resonant sites shifts towards the neutral line, since Alfvén resonance necessarily occurs where the phase velocity of the driver is equal to the local Alfvén speed, $\omega_{0}/k=v_{a}(x)$. When the resonant layers, which are located on either side of the neutral line, overlap across the neutral line as a consequence of their finite widths, the resonance couples with the reconnection mode. The subsequent section discusses the effect of finite frequency on such a coupling.
@@ VI.3 Discussion: scalings laws and effect of finite frequency on reconnection
Figure 5: Plots of the reconnection rate and reconnected flux as functions of $\omega_{0}$ evaluated at the time of maximum reconnection rate in the constant-psi regime, for $S=10^{8}$. They are calculated from the inverse Laplace transforms of Eq. (49), divided by $S$, and Eq. (44), which are applicable for $\omega_{0}\ll S^{-1/3}$. The red (left most) line in each plot indicates $\omega_{0}=S^{-3/5}$. The orange (middle) line indicates the location of half maximum half width of $\psi_{0}$. The green (right most) line indicates $\omega_{0}=S^{-1/3}.$ In the frequency range beyond the half-width of $\psi_{0}$ but below $\omega_{0}=S^{-1/3}$, we observe a power law scaling $R\sim\omega_{0}^{-0.25}$ and $\psi_{0}\sim\omega_{0}^{-1.23}$.
@@
To further understand the effect of finite frequency on forced reconnection, in this section we analyze two sets of scalings predicted by the theory. The first set is the scaling of the width $\delta$ of the boundary layer for forced reconnection and Alfvén resonance as functions of the Lundquist number $S$. The second set is the scaling of the constant-psi reconnection rate and reconnected flux at the neutral line as functions of the driving frequency $\omega_{0}$.
@@
Through inspection of the current density equation in the constant-psi regime which describes forced reconnection, Eq. (48), we find the theoretical scaling law for $\delta$. For this regime we assume the scaling $g\sim S^{-3/5}$, found from the poles in the denominator. Upon substitution of $g$ for $S$ in Eq. (48) we find the argument of the Modified Bessel and Struve functions to be of order unity when $x\sim S^{-2/5}$. This indicates a layer width scaling $\delta\sim S^{-2/5}$, which is the known scaling for this process [2].
@@
For Alfvén resonance we perform the same analysis of its equivalent current density equation, applicable for large driving frequencies $\omega_{0}\gg S^{-1/3}$, Eq. (54). For this scaling we need only the argument of the exponential to determine the inner layer width dependence on $S$. We see that the argument is of order unity when $\theta\sim S^{1/3}$ which is indicative of a layer width scaling $\delta\sim S^{-1/3}$, as expected [9, 8, 10].
@@
Although it is not possible to provide a theoretical solution for the intermediate regime when $Q\sim 1$ (which characterizes the region at which our approximations mismatch with simulation in Figure 4), we can nevertheless use the general solution given in Eq. (28) to infer the expected $\delta$. From the definition of the parameter $Q$, Eq. (23), the scaling $g\sim S^{-1/3}$ naturally emerges. Replacing $g\sim S^{-1/3}$ in Eq. (28) we find the argument of the Tricomi confluent hypergeometric functions of Eq. (28), $z=\sqrt{Q}\theta/(Sk)^{2/3}$, is of order 1 for $\theta\sim S^{1/3}$ or $\delta\sim S^{-1/3}$. The regime $Q\sim 1$, which corresponds to $t\sim S^{1/3}$, applies to both Alfvén resonance and forced reconnection. In particular, it shows that forced reconnection starts off in the non-constant-psi regime, as first discussed in [2], before continuing on through the constant-psi. Thus, we see that when resistivity becomes important at $t\sim S^{1/3}$ both forced reconnection and Alfvén resonance exhibit the same layer width, allowing for their coupling, when the forcing frequency is sufficiently small.
@@
We make now analytical predictions of how the reconnection rate and the reconnected flux scale with the forcing frequency $\omega_{0}$. We conduct this analysis under the constant-psi regime, i.e., $R$ is calculated by using Eq. (49) divided by $S$, and (44) is used for $\psi_{0}$, that in this approximation represents the reconnected flux at $x=0$.
@@
For frequencies $\omega_{0}\ll S^{-3/5}$, the system has time to evolve from the ideal to the constant-psi regime almost unaffected by the oscillating driver, and we expect that for such low frequencies both $R$ and $\psi_{0}$ are unaffected by $\omega_{0}$. When the frequency approaches the inverse of the constant-psi time scale, which is around the time when the peak in reconnection occurs (see Fig. 4), we expect the reconnection process to be affected by the finite frequency driver. Now, in the frequency range $S^{-3/5}\ll\omega_{0}\ll S^{-1/3}$ one can neglect the term $\Delta_{0}=2k_{0}/\tanh(k_{0})$ in the denominator of Eqs. (44) and (49). Additionally, as demonstrated previously, as $\omega_{0}$ increases the contribution to the long time solution is provided by residues at $g=\pm i\omega_{0}$. By analyzing under this light the residues of the inverse Laplace transforms of Eqs. (49) and (44) at $g=\pm i\omega_{0}$, it can be shown that the expected scalings are $R\sim\omega_{0}^{-1/4}$ and $\psi_{0}\sim\omega_{0}^{-5/4}$. This would suggest that although $\omega_{0}\sim S^{-1/3}$ determines the transition from forced reconnection to Alfvén resonance, the reconnection rate is reduced in the range of frequencies $S^{-3/5}\ll\omega_{0}\ll S^{-1/3}$. For frequencies larger than $\omega_{0}\gg S^{-1/3}$ reconnection is completely suppressed by the Alfvén resonance.
@@
To measure the effect of finite frequency on reconnection, in Figure 5 we show the plot of the normalized, maximum reconnection rate $R$ and corresponding reconnected flux function $\psi_{0}$ as functions of $\omega_{0}$ in the left and right panel, respectively, for $S=10^{8}$. As explained above, the constant-psi regime was assumed in making these plots. Therefore, we used Eqs. (49), divided by $S$, for $R$ and (44) for $\psi_{0}$, and calculated their inverse Laplace transforms at the time for which the reconnection rate is at its maximum. With the vertical red, orange and green (left, central, and right) lines, we mark $\omega_{0}=S^{-3/5}$, the half-width of $\psi_{0}$, and $\omega_{0}=S^{-1/3}$. As can be seen, the predicted scalings of $R\sim\omega_{0}^{-1/4}$ and $\psi_{0}\sim\omega_{0}^{-5/4}$ are very closely matched in Figure 5, in which we found power law scalings of $R\sim\omega_{0}^{-0.25}$ and $\psi_{0}\sim\omega_{0}^{-1.23}$ for the reconnection rate and reconnected flux, respectively, shown by the fits with the dashed line. That is we find, as expected, a strong suppressing effect on reconnection for frequencies less than the transition frequency $\omega_{0}\sim S^{-1/3}$ and higher than $\omega_{0}\sim S^{-3/5}$.